CASES99: Logbook Entries

CASES99: Site stn3 Messages, 5 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
163 Sat 23-Oct-1999Cleaning Kryptonsstn3delany
65 Sat 09-Oct-1999cleaned krypton on stn3stn3oncley
64 Sat 09-Oct-1999T/RH needed new fan stn3oncley
37 Wed 29-Sep-1999Krypton spiking problem w/EVE A/Dstn3militzer
36 Wed 29-Sep-1999Tsfc replaced on PAM-3stn3militzer


163: Kryptons, Site stn3, Sat 23-Oct-1999 11:53:45 CDT, Cleaning Kryptons
Between 1100 and 1145, the Krypton hygrometers at stations 2, 5, 4 and 3 
were cleaned using water wetted Q-tips and were then dried using Q-tips
65: Kryptons, Site stn3, Sat 09-Oct-1999 16:01:51 CDT, cleaned krypton on stn3
I just reached up to clean the krypton on stn3 about 10 min. ago.  It may
take a few minutes for this sensor to recover.

This done because Tom had noticed low fluxes on this sensor.
64: TRH, Site stn3, Sat 09-Oct-1999 15:59:51 CDT, T/RH needed new fan
The fan had died on stn3 T/RH.  An hour ago, I pulled the sensor off this
station, replaced the fan with a new one, and just reinstalled the sensor.

Tom thinks this has been bad since at least 29 Sep.


37: Kryptons, Site stn3, Wed 29-Sep-1999 16:01:29 CDT, Krypton spiking problem w/EVE A/D
Krypton data from sensor 1391 was glitching as shown on the
cockpit display.  It was the only sensor showing this trend.
A check running 'adflag' of the station revealed apparent
dropped a/d counts on the first cycle of the 20hz krypton task
after a reading taken of the .5hz power monitor parameters.
Approx 50 counts would be missed.
This is symptomatic of the a/d problem usually experienced
by the power monitor which can have a reduced vbatt reading.

Station 3 electronics box has been the poorest performing
PAM station in terms of responsiveness with the fastout and
high speed ingest processing.  A couple of vx errors were
observed.  This is probably a timing problem between the two
analog tasks and the device driver.   It is likely either
the a/d board, dcom332 cpu or the backplane or a combination.

A/D board swap to station 1
From station 3 board 439220 removed and swapped into sta 1
From station 1 board 439176 removed and put in sta 3.
This should reveal whether the glitching problem follows the
A/D board.  I don't think it will.
36: Radiation, Site stn3, Wed 29-Sep-1999 15:58:52 CDT, Tsfc replaced on PAM-3
Everest IRT Swapped:
Sensor 3577-4 removed; sensor 3577-2 installed

After the thunder/rain storm 2 nights ago the old sensor
was reading 100 deg-C.